Two Carnot engine operate in series. The first engine at 852c and reject to a reservoir at TC.the second engine receive this heat and reject it to a reservoir at 47c.it is observe that the efficiency of the engines are equal.find the value of t

Grade:12th pass

2 Answers

Arun
4 years ago
Dear student
 
efficiency 1 = efficiency 2
 
1 – T / 1125 = 1 – 320/T
 
320 /T =  T / 1125
 
T^2 = 320*1125 = 360000
 
hence T = 600 K
 
T = 600 – 273 = 327 degree
Khimraj
4 years ago

For Carnot Engine 1 T c = T a n d T h = 1200 K Efficiency {eq}\eta_1 \ = ( 1 \ - \frac{T_c}{T_h}) \ = ( 1 \ -...